OpenStudy (anonymous):

The degree measure of three angles of a triangle are x, y, and z. if z=x+y/2, then find the value of z

OpenStudy (ajprincess):

Sum of the interior angles of a triangle=180 That is x+y+z=180

OpenStudy (anonymous):

@ajprincess thats it????? .....hw is tht the answer???

OpenStudy (ajprincess):

That's nt the answer. What is z?

OpenStudy (ajprincess):

is it \[\frac{x+y}{2}\]?

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

z=(x+y)/2 multiply by 2 both sides. Then u will get the value of x+y in terms of z. Put it in x+y+z=180 and then solve for z. Can u do @basith?

OpenStudy (anonymous):

is it 60

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

yup it is

OpenStudy (anonymous):

thnx

OpenStudy (ajprincess):

yw:)